OpenStudy (anonymous):

Which are partial products for 56 × 31? Choose all answers that are correct. A. 6 B. 50 C. 150 D. 180

OpenStudy (anonymous):

pls answer

OpenStudy (anonymous):

so b and a

OpenStudy (anonymous):

hold up, no I was thinking of something else.

OpenStudy (anonymous):

cuz i need to know the direct answer

OpenStudy (anonymous):

56*31 = (50+ 6)(30 + 1) = (50* 30) + (50 * 1) + (6* 30) + (6 * 1)

OpenStudy (anonymous):

pls and thank you

OpenStudy (lolacole12):

Ok, so: 56 × 31 (50 + 6) (30 + 1) 50 x 30 50 x 1 6 x 30 6 x 1

OpenStudy (anonymous):

so the partial products are 1500, 50, 180, and 6

OpenStudy (zale101):

Lor_aka you are here to get walked through the answer, not for direct answer. We can help you to climb the ladder.

OpenStudy (anonymous):

ok thx for all the help